### v2

@@ Abstract
In this second paper on the variability survey of central stars of planetary nebulae (CSPNe) using ZTF, we focus on the 11 long-timescale variables with variability timescales ranging from months to years. We also present preliminary analyses based on spectroscopic and/or photometric follow-up observations for six of them. Among them is NGC 6833, which shows a day periodic variability with strange characteristics: ‘triangle-shaped’ brightening in , , and WISE bands but almost coincidental shallow dips in the -band. We speculate this to be a wide but eccentric binary with the same orbital period. Long-period near-sinusoidal variability was detected in two other systems, NGC 6905 and Kn 26, with periods of days and days, respectively, making them additional wide-binary candidates. The latter also shows a short period at hours which can either be from a close inner binary or pulsational origin. We present CTSS 2 and PN K 3-5 which show brightening and significant reddening over the whole ZTF baseline. A stellar model fit to the optical spectrum of CTSS 2 reveals it to be one of the youngest post-AGB CSPN known. Both show high-density emission-line cores. These appear to be late thermal pulse candidates, currently evolving towards the AGB phase, though alternative explanations are possible. We then present recent HST/COS ultraviolet spectroscopy of the known wide-binary candidate LoTr 1 showing that the hot star is a spectroscopic twin of the extremely hot white dwarf in UCAC2 46706450. We think that the long photometric period of 11 years is the binary orbital period. Finally, we briefly discuss the ZTF light curves of the remaining variables, namely Tan 2, K 3-20, WHTZ 3, Kn J1857+3931, and IPHAS J1927+0814. With these examples, we present the effectiveness of the von Neumann statistics and Pearson Skew-based metric space in searching for long-timescale variables.
@@ 1 Introduction
Over the past decade, studies of photometric variability of central stars of planetary nebulae (CSPNe) have significantly improved our understanding of these systems. The majority of past works have been dedicated to the search for binarity in CSPNe in support of the common-envelope (CE) formation channel of planetary nebulae (PNe; see, for example, Bond et al. 1978 and Bond & Livio 1990 for some of the first discoveries). Wide-field photometric surveys like the Optical Gravitational Lensing Experiment (OGLE), Kepler-K2, Transiting Exoplanet Survey Satellite (TESS), and recently the Zwicky Transient Facility (ZTF) have together discovered more than a hundred binary CSPNe (Miszalski et al., 2009a, b; Jacoby et al., 2021; Aller et al., 2020, 2024; Bhattacharjee et al., 2025; Chen et al., 2025). Periodicity has also led to the characterization of other phenomena like rotational variability caused by stellar spots (see for example Werner et al. 2019, 2023; Bond & Zeimann 2024), or pulsations of the hot CSPN (Bond & Meakes, 1990; Ciardullo & Bond, 1996; Hajduk et al., 2014; Córsico et al., 2021).
@@
Though much less explored compared to periodic variables, CSPNe also exhibit aperiodic variability. One of the causes is strong winds in CSPNe, mostly causing ‘jittering’ variability (see for example Handler et al. 1997; Arkhipova et al. 2012, 2013; Córsico et al. 2021). Unusual and large-amplitude aperiodic photometric variability of CSPNe has also led to the discoveries of R Coronae Borealis variables inside PNe, providing significant insights into the late thermal pulse formation channel of this class of stars (Gonzalez et al., 1998; Asplund et al., 1997; Clayton & De Marco, 1997; Jeffery & Hambsch, 2019; Rosenbush & Efimov, 2015). Recently, transits of dust and debris discs was proposed as an explanation for the deep quasi-periodic dips seen in the nucleus of the planetary nebula WeSb 1 (Bhattacharjee et al., 2025; Budaj et al., 2025). These demonstrate the variety of causes that can induce photometric variability in CSPNe.
@@
Most of the aforementioned phenomena occur on short timescales, ranging from a few minutes to a few days. On the other hand, there have only been very few instances of long-timescale photometric modulation in CSPNe. One example is PN IC 4997 (Aller & Liller, 1966) which shows decade-long brightening and dimming periods which may result from either a long-period binary (Kostyakova & Arkhipova, 2009) or wind/mass loss variations (Arkhipova et al., 2020). Long-period light curve modulation due to binarity is seen in LoTr 5 (period of 7.4 years, Kővári et al. 2019). Long-term photometric monitoring also resulted in the discovery of the day binary in pre-PN V510 Pup (Manick et al., 2021). Post-asymptotic giant branch (AGB) stars undergoing (very) late thermal pulses also show long-timescale photometric changes (see for example van Genderen & Gautschy 1995 for FG Sge and Schaefer & Edwards 2015 for Stingray Nebula). There are a few other similar examples in the literature. However, there has not been any systematic search for long-timescale variability in CSPNe.
@@
This is the second of a two-paper series conducting a systematic study of CSPN variability with Zwicky Transient Facility (ZTF, Graham et al. 2019; Masci et al. 2019). Here, we focus on the CSPNe showing long-timescale photometric variability in ZTF (the first one focused on the short-timescale variables, Bhattacharjee et al. 2025, hereafter Paper I). We find 11 objects showing prominent long-term variations in the light curve, making this by far the largest study of long-timescale variability in CSPNe. The rest of the paper is as follows. In section 2, we briefly review the methodology employed in selecting the variable sources. In section 3 we present our results. This includes spectroscopic observations of some of the objects and related analyses. Finally, we present our conclusions in section 4.
@@ 2 Method and Sample Selection
The detailed procedure for the selection of variable CSPNe in ZTF data has been described in Paper I. We briefly review the methodology here. Our starting sample is the Gaia-EDR3-selected list of CSPNe presented in Chornay & Walton (2021). We restrict ourselves to sources with CSPN identification reliability , resulting in 1812 CSPNe, of which ZTF data is recovered for 991 sources. Following this, we applied several photometric quality cuts, primarily to avoid sources where the extended nebula significantly affects the photometry. This reduced the number of sources to 490. To quantify variability, we calculate the Normalized Excess Variance metric (NEV) for the light curves. Using an appropriate cut on this metric value, we arrive at our final list of 94 highly variable CSPNe (HNEV sample). Among the HNEV sources, we classify 83 objects to be short-timescale variables (the focus of Paper I). The variability of the remaining 11 objects occurs on much longer timescales (months to years) and is well resolved with ZTF cadence. We classify these objects as long-timescale variables, which are the focus of this paper. The list of the 11 long-timescale variables is presented in Table 3.
In Paper I, we had presented a two-dimensional metric space defined by the metrics von Neumann statistic (vonN) and Pearson Skew (SkewP) as a simple and novel technique to identify “exotic variables”. Briefly, the former metric quantifies the orderliness in the variability (against random variability from photon noise), and the latter differentiates between dimming and brightening events. We discussed that objects with low vonN () are often “interesting”, i.e., objects which can teach us new phenomena. We demonstrated that the debris disk candidate WeSb 1 lies in the low vonN space and its large negative SkewP successfully represents the deep eclipses. In Appendix A, we identify the long-timescale variables in this space and show its effectiveness in identifying them from large data sets. Such a discovery tool is expected to come in handy with upcoming large datasets with the advent of new surveys like Large Synoptic Survey Telescope (LSST, Ivezić et al. 2019).
@@ 3 Results and Analyses
| PNG | Name | RA | Dec | GaiaID | PNStat | PNRad | Rel | Known Period | ZTF Period |
| 082.5+11.3 | NGC 6833 | 297.4440 | 48.9611 | 2086764737566091136 | T | 0.3 | 1.0 | – | 978 d |
@@
| K 3-20 | Long-timescale brightening (symmetric in appearence) associated with redenning. Part of a longer-periodic binary light curve? |
@@
Below we present some discussion on the long-timescale variables. The wide variety of light curves (and, thus, the underlying causes of variability) makes the detailed study of all the objects together difficult. We are in the process of obtaining follow-up spectroscopic and photometric observations of these objects. As a demonstration of this effort, we present the preliminary analyses and discussion on several of the objects based on the available data. Before diving into the details, we summarize our analyses in Table 2 for the convenience of the readers.
We note here that except WHTZ 3, the ZTF variability of all the other objects below were recently reported in a completely independent work by (Chen et al., 2025, henceforth, C25). Though they successfully present a broad range of discoveries from ZTF with much less stringent cuts than the ones applied here (going beyond the HNEV sample in our papers), they provide no additional discussion on the discovered systems. This makes this work different as we present follow-up observations and extensive discussions. These shall serve as the first step in understanding these complex objects. But we refer to C25 whenever needed in our discussions.
All of these objects appear in the catalog of PN (candidates) detected by GALEX (Gómez-Muñoz et al., 2023). Five among these objects (LoTr 1, Kn J1857+3931, NGC 6833, NGC 6905, and Kn 26) also appear in the catalog of hot subluminous stars by Culpan et al. (2022).
@@ 3.1 NGC 6833
NGC 6833 is a well-known PN and a True PN in HASH222Hong Kong/Australian Astronomical Observatory/Strasbourg Observatory H Planetary Nebula (HASH, Parker et al. 2016) catalog is the largest catalog of PNe which serves as an important reference for Paper I and this work. It ranks its PN candidate entries as “True”, “Likely” and “Probable” PNe, in the decreasing order of their likeliness of actually being a PN. The True PNe are spectroscopically verified to be so. with an apparent diameter of . Due to its compact angular size, current ground-based surveys cannot resolve the central star from the nebula. Thus, the recorded fluxes are a combination of both the star and the nebula. Images from the Hubble Space Telescope (see for example Hyung et al. 2010) reveal the box-like shape of the PN. NGC 6833 has found its presence in various catalogs over decades and has been the subject of several spectroscopic studies. The photometric behavior, however, remained unnoticed.

### v3

@@ Abstract
In this second paper on our variability survey of central stars of planetary nebulae (CSPNe) using the Zwicky Transient Facility (ZTF), we report 11 long-timescale variables with variability timescales ranging from months to years. We also present preliminary analyses based on spectroscopic and/or photometric follow-up observations for six of them. Among them is NGC 6833, which shows a day periodic variability with strange characteristics: ‘triangle-shaped’ brightening in , , and WISE bands but almost coincidental shallow dips in the -band. The most plausible explanation is a wide binary with the photometric period being the orbital period. Long-period near-sinusoidal variability was detected in two other systems, NGC 6905 and Kn 26, with periods of days and days, respectively, making them additional wide-binary candidates. The latter also shows a short period at hours. We then present CTSS 2 and K 3-5, which show brightening and significant reddening over the whole ZTF baseline. A stellar model fit to the optical spectrum of CTSS 2 reveals it to be one of the youngest post-AGB CSPNe known. Both show high-density emission-line cores. We propose these to be late-thermal-pulse candidates, currently evolving towards the AGB phase. We then present recent HST/COS ultraviolet spectroscopy of the known wide-binary candidate LoTr 1, showing that the hot star is a spectroscopic twin of the extremely hot white dwarf in UCAC2 46706450. Similar to this object, LoTr 1 also has a fast rotating wide subgiant companion. We suggest that the long photometric period of 11 years is the binary orbital period. Finally, we briefly discuss the ZTF light curves of the remaining variables, namely Tan 2, K 3-20, WHTZ 3, Kn J1857+3931, and IPHAS J1927+0814. With these examples, we present the effectiveness of the von Neumann statistics and Pearson Skew-based metric space in searching for long-timescale variables.
@@ 1 Introduction
Over the past decade, studies of photometric variability of central stars of planetary nebulae (CSPNe) have significantly improved our understanding of these systems. The majority of past works have been dedicated to the search for binarity in CSPNe in support of the common-envelope (CE) formation channel of planetary nebulae (PNe; see e.g.; Bond & Livio 1990; Balick & Frank 2002; De Marco 2009; Jones & Boffin 2017; Boffin & Jones 2019). Wide-field photometric surveys like the Optical Gravitational Lensing Experiment (OGLE), Kepler-K2, Transiting Exoplanet Survey Satellite (TESS), and recently the Zwicky Transient Facility (ZTF) have together discovered more than a hundred binary CSPNe (Miszalski et al., 2009a, b; Jacoby et al., 2021; Aller et al., 2020, 2024; Bhattacharjee et al., 2025; Chen et al., 2025). Periodicity has also led to the characterization of other phenomena like rotational variability caused by stellar spots (see for example Werner et al. 2019; Bond & Zeimann 2024), or pulsations of the hot CSPN (Bond & Meakes, 1990; Ciardullo & Bond, 1996; Hajduk et al., 2014; Córsico et al., 2021).
@@
Though much less explored compared to periodic variables, CSPNe also exhibit aperiodic variability. One of the causes is strong winds in CSPNe, mostly causing ‘jittering’ variability (see for example Handler et al. 1997; Arkhipova et al. 2012, 2013; Córsico et al. 2021). Unusual and large-amplitude aperiodic photometric variability of CSPNe has also led to the discoveries of R Coronae Borealis variables inside PNe, providing significant insights into the late thermal pulse formation channel of this class of stars (Gonzalez et al., 1998; Asplund et al., 1997; Clayton & De Marco, 1997; Jeffery & Hambsch, 2019; Rosenbush & Efimov, 2015; Jeffery & Hambsch, 2019). Recently, transits of dust and debris discs was proposed as an explanation for the deep quasi-periodic dips seen in the nucleus of the planetary nebula WeSb 1 (Bhattacharjee et al., 2025; Budaj et al., 2025). These demonstrate the variety of causes that can induce photometric variability in CSPNe.
@@
Most of the aforementioned phenomena occur on short timescales, ranging from a few minutes to a few days. On the other hand, there have only been very few instances of long-timescale photometric modulation in CSPNe. One example is PN IC 4997 (Aller & Liller, 1966) which shows decade-long brightening and dimming periods, which may result from either a long-period binary (Kostyakova & Arkhipova, 2009) or wind/mass loss variations (Arkhipova et al., 2020). Long-period light curve modulation due to binarity is seen in LoTr 5 (period of 7.4 years, Kővári et al. 2019). Long-term photometric monitoring also resulted in the discovery of the day binary in pre-PN V510 Pup (Manick et al., 2021). Post-asymptotic giant branch (AGB) stars undergoing (very) late thermal pulses also show long-timescale photometric changes (van Genderen & Gautschy 1995 for FG Sge and Schaefer & Edwards 2015 for Stingray Nebula). There are a few other similar examples in the literature. However, there has not been any systematic search for long-timescale variability in CSPNe.
@@
This is the second paper of the series conducting a systematic study of CSPN variability with Zwicky Transient Facility (ZTF, Graham et al. 2019; Masci et al. 2019). Here, we focus on the CSPNe showing long-timescale photometric variability in ZTF (the first one focused on the short-timescale variables, Bhattacharjee et al. 2025, hereafter Paper I). We find 11 objects showing prominent long-term variations in the light curve, making this by far the largest study of long-timescale variability in CSPNe. The rest of the paper is as follows. In section 2, we briefly review the methodology employed in selecting the variable sources. In section 3 we present our results. This includes spectroscopic observations of some of the objects and related analyses. Finally, we present our conclusions in section 4.
@@ 2 Method and Sample Selection
The detailed procedure for the selection of variable CSPNe in ZTF data has been described in Paper I. We briefly review the methodology here. Our starting sample is the Gaia-EDR3-selected list of CSPNe presented in Chornay & Walton (2021). We restrict ourselves to sources with CSPN identification reliability , resulting in 1812 CSPNe, of which ZTF data are recovered for 991 sources. Following this, we applied several photometric quality cuts, primarily to avoid sources where the extended nebula significantly affects the photometry. This reduced the number of sources to 490. To quantify variability, we calculate the Normalized Excess Variance metric (NEV) for the light curves. Using an appropriate cut on this metric value, we arrive at our final list of 94 highly variable CSPNe (HNEV sample). Among the HNEV sources, we classify 83 objects to be short-timescale variables (the focus of Paper I). The variability of the remaining 11 objects occurs on much longer timescales (months to years) and is well resolved with ZTF cadence. We classify these objects as long-timescale variables, which are the focus of this paper. The list of the 11 long-timescale variables is presented in Table 1. For convenience of readers, we also provide upfront the summary of our analyses in Table 2, the details being presented in the respective sections.
In Paper I, we had presented a two-dimensional metric space defined by the metrics von Neumann statistic (vonN) and Pearson Skew (SkewP) as a simple and novel technique to identify “exotic variables.” Briefly, the former metric quantifies the orderliness in the variability (against random variability from photon noise), and the latter differentiates between dimming and brightening events. We discussed that objects with low vonN () are often “interesting,” i.e., objects which can teach us new phenomena. We demonstrated that the debris-disk candidate WeSb 1 lies in the low vonN space and its large negative SkewP successfully represents the deep eclipses. In Appendix A, we identify the long-timescale variables in this space and show its effectiveness in identifying them from large data sets. Such a discovery tool is expected to come in handy with upcoming large datasets with the advent of new surveys like the Legacy Survey of Space and Time (LSST, Ivezić et al. 2019).
@@ 3 Results and Analyses
| PNG | Name | RA (deg.) | Dec (deg.) | Gaia DR3 ID | PNStat | PNRad (′′) | Rel | Known Period | ZTF Period |
| 082.5+11.3 | NGC 6833 | 297.4440 | 48.9611 | 2086764737566091136 | T | 0.3 | 1.0 | – | 980 d |
@@
Note. — Description of the last five column names. PNStat: the PN status as reported in HASH. T and P stand for True and Probable, respectively. PNRad: radius of the PN as reported in HASH. Rel: CSPN assignment reliability fraction from Chornay & Walton (2021). Known Period: previously reported photometric period. ZTF Period: period(s) detected in ZTF light curve.
@@
| K 3-20 | Long-timescale brightening (symmetric in appearance) associated with reddening. Part of a longer-periodic binary light curve? |
@@
Below we present the objects showing long-term photometric variability. We note up front that, except WHTZ 3, the ZTF variability of all the other objects below was recently reported in a completely independent work by (Chen et al., 2025, henceforth, C25). However, they provide no discussion on the origin of the variability or supplement their discovery with follow-up observations. In this work, we provide follow-up spectroscopic and photometric observations of six of the objects, shedding light on the nature of the objects and the origin of their variability. Additionally, we list the remaining five variables and urge further research by the community.
All the objects discussed here appear in the catalog of PN (candidates) detected by GALEX (Gómez-Muñoz et al., 2023). Five among these objects (LoTr 1, Kn J1857+3931, NGC 6833, NGC 6905, and Kn 26) also appear in the catalog of hot subluminous stars by Culpan et al. (2022).
@@ 3.1 NGC 6833
NGC 6833 is a well-known PN and a “True PN” in the HASH catalog,222Hong Kong/Australian Astronomical Observatory/Strasbourg Observatory H Planetary Nebula (HASH, Parker et al. 2016) catalog is the largest catalog of PNe, serving as an important reference for Paper I and this work. It ranks its PN candidate entries as “True,” “Likely,” and “Probable” PNe, in decreasing order of their likeliness of actually being a PN. The True PNe are spectroscopically verified to be so. with an apparent diameter of . Due to its compact angular size, current ground-based surveys cannot resolve the central star from the nebula. Thus, the recorded fluxes are a combination of both the star and the nebula. Images from the Hubble Space Telescope (see for example Hyung et al. 2010) reveal a box-like shape of the PN. NGC 6833 is present in various catalogs over decades, and has been the subject of several spectroscopic studies. Its photometric variability, however, remained unnoticed.
3.1.1 Light Curve and Spectroscopy
The ZTF light curve for NGC 6833 is shown in Figure 1 (top panel). Fortunately, this target has good -band coverage. The most prominent feature is the regular brightening events in the ZTF and the redder bands, with symmetric ingress and egress. The amplitude is in the range of , with the largest amplitude in -band. We note here though that the nebular flux can reduce the perceived variability, so the intrinsic amplitude of the source (likely the central star) can be much higher (see a brief discussion in Section 3.1.2). The less prominent but more intriguing feature appears in the ZTF band, which exhibits shallow (amplitude of ) dips nearly coincident with the peaks in the redder bands. The behavior is periodic and significant peaks appear in the Lomb-Scargle periodogram333We first use the astropy.timeseries implementation of the Lomb-Scargle periodogram (Lomb, 1976; Scargle, 1982; VanderPlas, 2018) in all the three ZTF bands. The result is shown in the top panel of Figure 2. The detected periods in the three bands are marginally different (which is not surprising with long periods), with a mean value of days. The phase-folded ZTF light curve at this mean period is shown in the middle panel of the same figure. Here, the activity appears to be slightly phase-delayed from the redder to the bluer bands. However, the current data are insufficient to establish this property with high confidence.
Before any further discussion, we caution that the brightness of this object ( mag in and mag in ) is very close to the saturation limit of ZTF ( in all bands). In fact, in one of the ZTF fields, the reference image is indeed saturated in the -band (where the object is brightest, mainly due to the very strong [O III] emission lines). Though such data points get rejected through the imposed quality cuts, the chance remains that the other flux readings are unreliable. To ascertain that this has not significantly affected the observed variability in the -band, especially the dip, we checked the -band light curve from the All Sky Automated Survey for SuperNovae (ASAS-SN) light curve provided at Sky Patrol V2.0444http://asas-sn.ifa.hawaii.edu/skypatrol/ (Shappee et al. 2014; Hart et al. 2023), which has a larger saturation brightness ( magnitude). The ASAS-SN -band light curve displays similar behavior as ZTF and the dips are visible. The data quality, however, is worse than ZTF thus not shown. For ZTF and bands, manual inspection shows that neither the reference nor the science images are saturated. We thus proceed with the assumption that the light curves are trustworthy.
At the time of this work, the object was at its photometric extremum. Assuming the ZTF period to be stable, the best estimate places the previous optical observations of Wright et al. (2005) and Hyung et al. (2010) at the quiescent phase. This prompted us to obtain another spectrum, with the DouBle SPectrograph (DBSP; Oke & Gunn 1982) attached to the Cassegrain focus of the Palomar 200-in Hale Telescope, to look for any significant changes. However, we could not find any notable difference between the spectra. The emission-line fluxes also agree well within error bars. The new spectrum is provided in Appendix B.
We use the spectrum to determine if the observed photometric variability can arise from solely the variation in the emission lines. We consider the ZTF- band () which shows the largest amplitude of . We calculate the emission-line fluxes by subtracting the continuum. The emission line (including the Paschen series) contribution to the total flux is . With this spectrum taken at the -band maximum, this would require a significant decrement in the line strengths of all the emission lines to contribute meaningfully to the variability, which is unlikely. This leads us to conclude that there is a significant variation in the continuum. However, given the compact nature of the PN, it is not possible to separate the stellar and nebular continuum with the current data.
3.1.2 Discussion
The stable periodicity in the light curve strongly suggests binarity. The shape of the light curve with larger amplitude in redder bands555The strong emission lines, however, can significantly suppress any underlying continuum variability. To estimate the effect, we use the spectrum to subtract the emission lines and re-calculate the variability amplitudes in the ZTF bands. The amplitudes increased significantly. In fact, the ZTF band now has a higher amplitude than the band. This shows that the effect is significant and interpretation need to be made with sufficient care. suggests an irradiation effect. This would mean that the photometric period is the binary orbital period. This would make it the fourth-longest known binary period inside a PN (after BD+33 2642 Van Winckel et al. 2014; Jones et al. 2017, LoTr 5, and NGC 1514 Jones et al. 2017). The large amplitude at such a long period suggests an eccentric binary. This is not a surprise as the other known long-period binaries are also eccentric. Irradiation, however, is not the only possible explanation. The light curves in the red bands resemble several RV Tauri systems showing long-term photometric modulation (known as the RVb phenomenon, see for example the light curves in Kiss & Bódi 2017). In this case, the widely accepted explanation to the variability is the phase-dependent dust-obscuration of the flux-dominant star in the presence of a thick circumbinary disc (Van Winckel et al., 1999; Gezer et al., 2015). In this case, too, the photometric period represents the orbital period.
Neither of the two scenarios above can individually explain the anti-correlated behavior in the and the redder bands. A combination of both effects, however, could explain the variability. In an eccentric orbit, interaction (either directly between the stars or circumstellar disks) is enhanced near periastron. This can heat up the circumstellar dust leading to brightening in the red bands. Additionally, enhanced accretion can also lead to variability in emission lines, especially H. Such phenomena are sometimes observed during ”pulsed accretion” processes in protostars (see for example Muzerolle et al. 2013, though the corresponding period is much shorter) and AGB binaries (Chen et al., 2017; Bollen et al., 2017). Irradiation of the companion by the hot CSPN can also contribute to the brightening in the red bands. The dip in the blue band can then be caused by the heavy dust obscuration of the white dwarf, which is the dominant contributor in this wavelength range. The stochastic nature of the interaction may also explain any offset between the blue and red extrema.
We now consider possible cases where the photometric period does not relate to binarity. Rotational modulation is unlikely as, in most cases, the light curves in different bands are in phase. C25 discusses Canum Venaticorum (ACV, a group of magnetic and chemically peculiar stars) a possibility as they also show anti-correlation between blue and red bands (see for example Gröbel et al. 2017; Faltová et al. 2021). However, the periods are much shorter (all known objects having days, Faltová et al. 2021) compared to NGC 6833. We also consider the possibility of it being Be stars losing mass periodically due to pulsational instabilities. In most such cases however, the ingress time is significantly shorter than the egress time and the infrared amplitudes are much larger than observed in NGC 6833 (see, for example, the light curves presented in Froebrich et al. 2023). This makes this scenario, too, unlikely.
3.2 NGC 6905 and Kn 26
NGC 6905 is a spectroscopically well-studied PN. Its central star is a [WR]-type star with a temperature of 150 kK (Gómez-González et al., 2022). Spectral classifications of [WO] and [WC] have been given in the literature (e.g., Gómez-González et al. 2022 and Córsico et al. 2021). The star also shows pulsations with periods in the range of 700–900 seconds (Ciardullo & Bond, 1996; Córsico et al., 2021). The nebula shows a bipolar morphology.
Kn 26, on the other hand, is a known quadrupolar PN (Guerrero et al., 2013), with Lanning 384 being its central star. Guerrero et al. (2013) finds the SED to be consistent with a blackbody of kK. Unlike NGC 6905, Kn 26 has not been spectroscopically classified before. We obtained follow-up optical spectroscopy with the Large Binocular Telescope (LBT) which reveals it to have a weak emission-line type ([WELS]) nucleus (see Appendix C).